# Quality-aware Masked Diffusion Transformer for Enhanced Music Generation

### v1

@@ Abstract
In recent years, diffusion-based text-to-music (TTM) generation has gained prominence, *Equal contribution.
Code on https://github.com/ivcylc/qa-mdt
offering a novel approach to synthesizing musical content from textual descriptions. Achieving high accuracy and diversity in this generation process requires extensive, high-quality data, which often constitutes only a fraction of available datasets. Within open-source datasets, the prevalence of issues like mislabeling, weak labeling, unlabeled data, and low-quality music waveform significantly hampers the development of music generation models. To overcome these challenges, we introduce a novel quality-aware masked diffusion transformer (QA-MDT) approach that enables generative models to discern the quality of input music waveform during training. Building on the unique properties of musical signals, we have adapted and implemented a MDT model for TTM task, while further unveiling its distinct capacity for quality control. Moreover, we address the issue of low-quality captions with a caption refinement data processing approach. Our demo page is shown in https://qa-mdt.github.io/.
@@ 1 Introduction
Figure 1: The distribution curves of CLAP similarity and pseudo-MOS for large-scale open-source music databases AudioSet (Defferrard et al., 2016) and FMA (Defferrard et al., 2016), where darker areas represent higher text-audio alignment or audio quality.
@@
Text-to-Music (TTM) generation aims to transform textual descriptions of emotions, style, instruments, rhythm, and other aspects into corresponding music segments. This provides new expressive forms and innovative tools for multimedia creation. According to the scaling law principles (Peebles and Xie, 2023; Li et al., 2024), effective generative models require a large volume of training data. However, unlike previous image generation tasks  (Chen et al., 2024a; Rombach et al., 2021), acquiring high-quality music data often faces greater challenges. This is primarily due to copyright issues and the need for professional hardware equipment to capture high-quality music data. This makes building a high-performance TTM model particularly difficult.
@@
In the TTM field, high-quality paired data of text and music signals is scarce, this prevalent issue of low-quality data, highlighted in Figure 1, manifests in two primary challenges. Firstly, there is a weak correlation between music signals and captions, characterized by missing, weak, or incorrect captions, leading to low text-audio similarity which can be indicated by CLAP score (Wu* et al., 2023). Secondly, most available music signals often suffer from distortion due to noise, low recording quality, or outdated recordings, resulting in diminished generated quality which can be measured by pseudo-MOS scores from quality assessment models (Ragano et al., 2023). These challenges significantly hinder the training of high-performance music generation models, leading to poor rhythm, noise, and inconsistencies with textual control conditions in the generated audio. Therefore, effectively training on large-scale datasets with label mismatches, missing labels, or low-quality waveform has become an urgent issue to address.
@@
In this study, we introduce a novel quality-aware masked diffusion transformer (QA-MDT) to enhance music generation. This model effectively leverages extensive, open-source music databases, often containing data of varying quality, to produce high-quality and diverse music. During training, we inject music quality through quantified pseudo-MOS (p-MOS) scores into the denoising stage at multiple granularities to foster quality awareness, while coarse-level quality information is seamlessly integrated into the text encoder and fine-level details are embedded into the transformer based diffusion architecture. Masking strategy is also employed to enhance the spatial correlation of the music spectrum and further accelerates convergence. This innovative strategy guides the model in the generation phase to produce high-quality music by leveraging information corresponding to elevated p-MOS scores. Additionally, we utilize large language models (LLMs) and CLAP to synchronize music signals with captions, thereby enhancing text-audio correlation in extensive music datasets. Our ablation studies on public datasets confirm the effectiveness of our methodology, with the final model surpassing previous works in both objective and subjective measures.
@@ 2 Related Work
Text to music generation. Text-to-music generation aims to create music clips corresponding to input descriptive or summary text. Previous efforts have utilized either language models (LMs) or DMs to model quantized waveform representations or spectral features. For continuous waveform generation, MusicLM (Agostinelli et al., 2023), MusicGen (Copet et al., 2024), and MeLoDy (Lam et al., 2024) leverage LMs and DMs on residual codebooks obtained via quantization-based audio codecs (Zeghidour et al., 2021; Défossez et al., 2022). Moûsai (Schneider et al., 2023), Noise2Music (Huang et al., 2023), Riffusion (Forsgren and Martiros, 2022), and AudioLDM 2 (Liu et al., 2023a) use U-Net-related diffusion to model mel-spectrograms or latent representations obtained through a pretrained VAE, which are then converted into audio waveform via pretrained vocoders  (Kong et al., 2020). Existing methods lack consideration of low-quality data issues during the modeling phase and fail to effectively learn the distribution of high-quality audio. Our approach addresses this issue by introducing quality-aware training.
@@
Transformer based diffusion models. Traditional diffusion models typically use U-Net as the backbone, where the inductive biases of CNNs do not effectively model the spatial correlations of signals and are insensitive to scaling laws (Li et al., 2024). However, these issues have been effectively addressed in transformer-based diffusion models (Peebles and Xie, 2023). This advantage has been particularly evident in fields such as video generation (Brooks et al., ), image generation (Peebles and Xie, 2023; Chen et al., 2024a; Bao et al., 2022), and speech generation (Liu et al., 2023b). To expedite training and foster inter-domain learning of correlations, the masking strategy has been proven effective (Lei et al., ), which has yielded the best class-conditioned results on ImageNet (Gao et al., 2023), while a simpler architecture (Zheng et al., 2023) incorporating reconstruction losses and unmasked fine-tuning further enhances model training speed. However, these models have not yet been verified for text-controlled audio generation capabilities on large-scale audio datasets, and their adaptability to additional control information remains an open question.
@@ 3 Preliminary
Latent diffusion model. Direct application of DMs to cope with distributions of raw signals incurs significant computational overhead Ho et al. (2020); Song et al. (2020). Conversely, studies Liu et al. (2023c, a) apply them in a latent space with fewer dimensions. The latent representation $z_{0}$ is the ultimate prediction target for DMs, which involve two key processes: diffusion and reverse processes. In the diffusion process, Gaussian noise is incrementally added to the original representation at each time step, described by $z_{t+1}=\sqrt{1-\beta_{t}}z_{t}+\sqrt{\beta_{t}}\epsilon$, where $\epsilon$ is drawn from a standard normal distribution $\mathcal{N}(0,I)$, and $\beta_{t}$ is gradually adapted based on a preset schedule to progressively introduce noise into the state $z_{t}$. The cost function Ho et al. (2020); Liu et al. (2023c) is formalized as:
$\arg\min_{\theta}\left[\mathbb{E}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(z_{0},y),\epsilon}}\left[\left\|\epsilon-D_{\theta}\left(\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,t,y\right)\right\|^{2}\right]\right],$ (1)
where $D_{\theta}$, the denoising model, strives to estimate the Gaussian noise $\epsilon$, conditioned on the latent state $z_{t}$, the time step $t$, the conditional embedding $y$, and where $\alpha_{t}$ represents a predefined monotonically increasing function. In the reverse process, we obtain $z_{t-1}$ via the recursive equation: $z_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(z_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}\right)+\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}}\epsilon$, where $\epsilon_{\theta}$ represents the estimated Gaussian noise.
Classifier-free guidance and negative guidance. Classifier-free guidance (CFG), introduced by (Ho et al., 2020), increases the versatility of DMs by enabling both conditional and unconditional generation. Typically, a diffusion model generates content based on specific control signals $y$ within its denoising function $D_{\theta}(z_{t},t,y)$. CFG enhances this mechanism by incorporating an unconditional mode $D_{\theta}(z_{t},t,\emptyset)$, where $\emptyset$ symbolizes the absence of specific control signals. The CFG-enhanced denoising function is then expressed as $D_{\theta}^{\text{CFG}}(z_{t},t,y)=D_{\theta}(z_{t},t,y)+w(D_{\theta}(z_{t},t,y)-D_{\theta}(z_{t},t,\emptyset))$, where $w\geq 1$ denotes the guidance scale. During training, the model substitutes $y$ with $\emptyset$ at a constant probability $p_{\text{uncond}}$. In inference, $y$ might be replaced by a negative prompt like "low quality" to prevent the model from producing such attributes (Liu et al., 2023a).
@@ 4 Method
Method overview. We adopt a three-tiered strategy to enable the proposed model to predict high-quality music waveform across a complex corpus. Initially, we introduce a waveform-quality-aware technique in Section 4.1 that enables the model to discern waveform quality, extracted from from both coarse and fine-grain perspectives. Next, in Section 4.2, we present our mask diffusion model to enhance the model’s capacity for quality waveform modeling and the management of conditional text information. Lastly, we outline a novel approach to address issues with problematic text, such as weak, inaccurate, or missing captions.
4.1 Quality information injection
At the heart of our work lies the implementation of a pseudo-MOS scoring model (Ragano et al., 2023) to meticulously assign music quality to quality prefixes and quality tokens.
We define our training set as $\mathcal{D}_{o}=\{(M_{i},T^{o}_{i})\mid i=1,2,\dots,N_{D}\}$, where each $M_{i}$ represents a music signal and $T^{o}_{i}$ is the corresponding original textual signal, we achieve the p-MOS set $S=\{s_{1},s_{2},\ldots,s_{n}\}$. To optimize model learning from datasets with diverse audio quality and minimize the impact of low-quality audio, we initially assign p-MOS scores to each music track using a model fine-tuned with wav2vec 2.0 (Baevski et al., 2020) on a dataset of vinyl recordings for audio quality assessment. These scores facilitate dual-perspective quality control for enhanced granularity and precision.
First, We analyze this p-MOS set $S$ to identify a negative skew normal distribution with mean $\mu$ and variance $\sigma^{2}$. We define text prefixes based on $s$ as follows: prepend "low quality" if $s<\mu-2\sigma$, "medium quality" if $\mu-\sigma\leq m\leq\mu+\sigma$, and "high quality" if $s>\mu+2\sigma$. This information is prepended before processing through the text encoder with cross-attention, enabling the initial separation of quality-related information.
To achieve a more precise awareness and control of waveform quality, we synergize the role of text control with quality embedding. We define the quantization function $Q:[0,5]\to\{1,2,3,4,5\}$, which maps p-MOS scores to discrete quality levels based on the distance from the mean $\mu$ in terms of standard deviation $\sigma$:
@@ 4.1 Quality information injection
4.2 Quality-aware mask diffusion transformer
In a general patchify phrase with patch size $p_{f}\times p_{l}$ and overlap size $o_{f}\times o_{l}$, patchified token sequence $X=\{x_{1},x_{2},\dots,x_{P}\}\subset\mathbb{R}^{p_{f}\times p_{l}}$ are obtained through spliting the music latent space $\mathcal{M}_{spec}\in\mathbb{R}^{F\times L}$, as described in  5.2. The total number of patches $P$ is given by:
@@ 4.2 Quality-aware mask diffusion transformer
A 2D-Rope position embedding (Su et al., 2024) is added to each patch for better modeling of relative position relationship while a binary mask $\boldsymbol{m}\in\{0,1\}^{P}$ is applied during the training stage, with a variable mask ratio $\gamma$. This results in a subset of $\lfloor\gamma P\rfloor$ patches being masked that $\sum_{i=1}^{P_{N}}m_{i}=\lfloor\gamma P\rfloor$, leaving $P-\lfloor\gamma P\rfloor$ patches unmasked. The subset of masked tokens is invisible in the encoder stage and replaced with trainable mask tokens in the decoder stage following the same strategy utilized in AudioMAE (Huang et al., 2022) and MDT (Gao et al., 2023).
The transformer we use consists of $N$ encoder blocks, $M$ decoder blocks, and an intermediate layer to replace the masked part with trainable parameters. We treat the embedding of the quantized p-MOS score as a prefix token, concatenated with each stage’s music tokens. Let ${X}^{k}=[{x}^{k}_{1},{x}^{k}_{2},\ldots,{x}^{k}_{P}]\in\mathbb{R}^{P\times d}$ represent the output of $k$-th encoder or decoder block, where the initial input of the encoder $X^{0}=z_{t}=\alpha_{t}{z}_{0}+\sqrt{1-\alpha_{t}}\epsilon$, and the final decoder block estimate $X^{N+M}=z_{0}={X}^{0}=[{x}_{1},{x}_{2},\ldots,{x}_{P}]$. For $k<N$, indicating the encoder blocks, the sequence transformation focuses only on unmasked tokens:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{(z_{0},q_{vq},y),\epsilon}\left[\left\|\epsilon-D_{\theta}\left(\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,t,q_{vq},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y}\right)\right\|^{2}\right]$ (7)
@@
$D_{\theta}^{\text{High Quality}}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{high}}},y)=D_{\theta}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{high}}},y)+w(D_{\theta}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{high}}},y)-D_{\theta}(z_{t},t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{vq}^{\text{low}}},\emptyset))$ (8)
@@
4.3 Music caption refinement
@@
Initially, pretrained music caption model  (Doh et al., 2023) is employed to re-annotate each music signal, shown as $\mathcal{D}_{g}=\{(M_{i},T^{g}_{i})\mid T^{g}_{i}=\mathcal{F}_{\text{cap}}(M_{i}),i=1,2,\dots,N\}$. CLAP text-audio similarity is applied to filter $\mathcal{D}_{g}$ with a threshold of $\rho_{1}$, resulting in
@@
In this context, we meticulously filter out generated captions that do not correspond with their respective audio files. This misalignment may be attributed to inaccuracies within the captioner’s insufficient training. For the filtered data, we opt to retain the use of the original captions.
To ensure that valuable information from the original captions is not overlooked when using only the generated captions, we first need to filter out original captions suitable for the fusing stage. Formulated as
@@
The issue can stem from the original data being improperly labeled with terms such as ’speech, car’ from datasets like AudioSet (Gemmeke et al., 2017) and also may be because of desperately missing of labels.
@@
$\mathcal{D}_{\text{merge}}=\left\{(M_{i},T_{\text{fusion}})\mid\mathcal{S}(T^{o},T^{g})<\rho_{3},(M_{i},T^{o})\in\mathcal{D}^{\text{filter}}_{o},(M_{i},T^{g})\in\mathcal{D}^{\text{filter}}_{g}\right\}.$ (11)

### v2

# QA-MDT: Quality-aware Masked Diffusion Transformer for Enhanced Music Generation

@@ Abstract
In recent years, diffusion-based text-to-music (TTM) generation has gained prominence, offering an innovative approach to synthesizing musical content from textual descriptions. Achieving high accuracy and diversity in this generation process requires extensive, high-quality data, including both high-fidelity audio waveforms and detailed text descriptions, which often constitute only a small portion of available datasets. In open-source datasets, issues such as low-quality music waveforms, mislabeling, weak labeling, and unlabeled data significantly hinder the development of music generation models. To address these challenges, we propose a novel paradigm for high-quality music generation that incorporates a quality-aware training strategy, enabling generative models to discern the quality of input music waveforms during training. Leveraging the unique properties of musical signals, we first adapted and implemented a masked diffusion transformer (MDT) model for the TTM task, demonstrating its distinct capacity for quality control and enhanced musicality. Additionally, we address the issue of low-quality captions in TTM with a caption refinement data processing approach. Experiments demonstrate our state-of-the-art (SOTA) performance on MusicCaps and the Song-Describer Dataset. Our demo page can be accessed at https://qa-mdt.github.io/.
@@ 1 Introduction
footnotetext: * Equal Contribution
code and checkpoints on https://github.com/ivcylc/qa-mdt
@@
Text-to-music (TTM) generation aims to transform textual descriptions of emotions, style, instruments, rhythm, and other aspects into corresponding music segments, providing new expressive forms and innovative tools for multimedia creation. According to scaling law principles (Peebles and Xie 2023; Li et al. 2024a), effective generative models require a large volume of training data. However, unlike image generation tasks (Chen et al. 2024a; Rombach et al. 2021), acquiring high-quality music data often presents greater challenges, primarily due to copyright issues and the need for professional hardware to capture high-quality music. These factors make building a high-performance TTM model particularly difficult.
@@
In the TTM field, high-quality paired data of text and music signals is scarce. This prevalent issue of low-quality data, highlighted in Figure 1, manifests in two primary challenges. Firstly, most available music signals often suffer from distortion due to noise, low recording quality, or outdated recordings, resulting in diminished generated quality, as measured by pseudo-MOS scores from quality assessment models (Ragano, Benetos, and Hines 2023). Secondly, there is a weak correlation between music signals and captions, characterized by missing, weak, or incorrect captions, leading to low text-audio similarity, which can be indicated by CLAP scores (Wu* et al. 2023). These challenges significantly hinder the training of high-performance music generation models, resulting in poor rhythm, noise, and inconsistencies with textual control conditions in the generated audio. Therefore, effectively training on large-scale datasets with label mismatches, missing labels, or low-quality waveforms has become an urgent issue to address.
Figure 1: The distribution curves of CLAP similarity and pseudo-MOS for large-scale open-source music databases AudioSet (Defferrard et al. 2016) and FMA (Defferrard et al. 2016), where darker areas represent higher text-audio alignment or audio quality.
@@
In this study, we introduce a novel quality-aware masked diffusion transformer (QA-MDT) to enhance music generation. This model effectively leverages extensive, open-source music databases, often containing data of varying quality, to produce high-quality and diverse music. During training, we inject quantified music pseudo-MOS (p-MOS) scores into the denoising stage at multiple granularities to foster quality awareness, with coarse-level quality information seamlessly integrated into the text encoder and fine-level details embedded into the transformer-based diffusion architecture. A masking strategy is also employed to enhance the spatial correlation of the music spectrum and further accelerate convergence. This innovative approach guides the model during the generation phase to produce high-quality music by leveraging information associated with elevated p-MOS scores. Additionally, we utilize large language models (LLMs) and CLAP model to synchronize music signals with captions, thereby enhancing text-audio correlation in extensive music datasets. Our ablation studies on public datasets confirm the effectiveness of our methodology, with the final model surpassing previous works in both objective and subjective measures. The main contributions of this study are as follows:
  • We propose a quality-aware training paradigm that enables the model to perceive the quality of the dataset during training, thereby achieving superior music generation in terms of both musicality and audio quality.
  • We innovatively introduced the Masked Diffusion Transformer to music signals, demonstrating its unique efficacy in modeling music latent space and its exceptional capability in perceiving quality control, thereby further improving both the generated quality and musicality.
  • We address the issue of low text-audio correlation in large-scale music datasets for TTM, effectively improving text alignment and generative diversity.
@@ 2 Related Work
Text-to-music generation aims to create music clips that correspond to input descriptive or summary text. Previous efforts have utilized either language models (LMs) or diffusion models (DMs) to model quantized waveform representations or spectral features. Models like MusicLM (Agostinelli et al. 2023), MusicGen (Copet et al. 2024), MeLoDy (Lam et al. 2024), and Jen-1 (Li et al. 2024b) leverage LMs and DMs on residual codebooks obtained via quantization-based codecs (Zeghidour et al. 2021; Défossez et al. 2022). Moûsai (Schneider et al. 2023), Noise2Music (Huang et al. 2023a), Riffusion (Forsgren and Martiros 2022), AudioLDM 2 (Liu et al. 2023b), and Stable Audio (Evans et al. 2024a) use U-Net-related diffusion to model mel-spectrograms or latent representations obtained through compression networks. Although some approaches attempt to guide the model towards generating high-quality content by setting negative prompts like “low quality” (Liu et al. 2023b; Chen et al. 2024b), few explicitly inject quality information during training. This results in the model’s inability to effectively perceive and control content quality.
Transformer based diffusion models.
@@
Traditional diffusion models typically use U-Net as the backbone, where the inductive biases of CNNs do not effectively model the spatial correlations of signals and are insensitive to scaling laws (Li et al. 2024a). However, transformer-based diffusion models (DiT) (Peebles and Xie 2023) have effectively addressed these issues. This advantage is particularly evident in fields such as video generation (Brooks et al. ), image generation (Peebles and Xie 2023; Chen et al. 2024a; Bao et al. 2022), and speech generation (Liu et al. 2023a). To expedite training and foster inter-domain learning of correlations, the masking strategy has proven effective, yielding SOTA class-conditioned performances on ImageNet (Gao et al. 2023). Additionally, a simpler architecture (Zheng et al. 2023) incorporating reconstruction losses and unmasked fine-tuning further enhances model training speed. However, these models have not yet been verified for text-controlled music generation on large-scale music datasets, and their adaptability with additional control information remains an open question. Make-an-audio 2 (Huang et al. 2023b) and, more recently, Stable Audio 2 (Evans et al. 2024b), have explored the DiT architecture for audio and sound generation. However, their approach models latent tokens by segmenting only along the time dimension to control and extend generation duration. In contrast, our focus is on finer segmentation within the latent space across both time and frequency, aiming for more precise modeling of music signals.
Quality enhancement in audio domain.
Previous research has made efforts to improve the quality of generated audio, particularly in two key areas: waveform fidelity and the alignment between input text and generated content. Waveform quality can be compromised by issues like aliasing from low sampling rates and limited expressiveness due to monophonic representations, while models like MusicGen (Copet et al. 2024) and Stable Audio (Evans et al. 2024a, b), which directly model 32k and 44.1k stereo audio, have significantly enhanced perceptual quality. Despite higher sampling rates and channels, the quality of audio in training datasets remains inconsistent, often suffering from noise, dullness, and a lack of rhythm or structure. These problems, often reflected by the Mean Opinion Score (MOS), are rarely addressed. In terms of text-audio alignment, Make-an-audio 2 (Huang et al. 2023b) and WavCaps (Mei et al. 2024) have employed ChatGPT-assisted data augmentation to improve temporal relationships and accuracy in audio effect generation. Although studies like Music-llama (Liu et al. 2024) and LP-musiccaps (Doh et al. 2023) have introduced captioning approaches for music, few have explored the augmentation and utilization of synthetic data in large-scale music generation tasks.
@@ 3 Preliminary
Latent diffusion model.
Direct application of DMs to cope with distributions of raw signals incurs significant computational overhead (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020). Conversely, studies (Liu et al. 2023c, b) apply them in a latent space with fewer dimensions. The latent representation $z_{0}$ is the ultimate prediction target for DMs, which involve two key processes: diffusion and reverse processes. In the diffusion process, Gaussian noise is incrementally added to the original representation at each time step $t$, described by $z_{t+1}=\sqrt{1-\beta_{t}}z_{t}+\sqrt{\beta_{t}}\epsilon$, where $\epsilon$ is drawn from a standard normal distribution $\mathcal{N}(0,I)$, and $\beta_{t}$ is gradually adapted based on a preset schedule to progressively introduce noise into the state $z_{t}$. The cost function (Ho, Jain, and Abbeel 2020; Liu et al. 2023c) is formalized as $\arg\min_{\theta}\mathbb{E}_{{(z_{0},y),\epsilon}}\left[\left\|\epsilon-D_{\theta}\left(\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,t,y\right)\right\|^{2}\right]$. where $D_{\theta}$, the denoising model, strives to estimate the Gaussian noise $\epsilon$, conditioned on the latent state $z_{t}$, the time step $t$, the conditional embedding $y$, and where $\alpha_{t}$ represents a predefined monotonically increasing function. In the reverse process, we obtain $z_{t-1}$ via the recursive equation: $z_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(z_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}\right)+\sqrt{\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}}\epsilon$, where $\epsilon_{\theta}$ represents the estimated Gaussian noise.
Classifier-free guidance.
Classifier-free guidance (CFG), introduced by (Ho, Jain, and Abbeel 2020), increases the versatility of DMs by enabling both conditional and unconditional generation. Typically, a diffusion model generates content based on specific control signals $y$ within its denoising function $D_{\theta}(z_{t},t,y)$. CFG enhances this mechanism by incorporating an unconditional mode $D_{\theta}(z_{t},t,\emptyset)$, where $\emptyset$ symbolizes the absence of specific control signals. The CFG-enhanced denoising function is then expressed as $D_{\theta}^{\text{CFG}}(z_{t},t,y)=D_{\theta}(z_{t},t,y)+w(D_{\theta}(z_{t},t,y)-D_{\theta}(z_{t},t,\emptyset))$, where $w\geq 1$ denotes the guidance scale. During training, the model substitutes $y$ with $\emptyset$ at a constant probability $p_{\text{uncond}}$. In inference, $\emptyset$ might be replaced by a negative prompt like “low quality” to prevent the model from producing such attributes (Liu et al. 2023b).
@@ 4 Method
Quality Information Injection
At the heart of our work lies the implementation of a pseudo-MOS scoring model (Ragano, Benetos, and Hines 2023) to meticulously assign music quality to quality prefixes and quality tokens.
We define our training set as $\mathcal{D}_{o}=\{(M_{i},T^{o}_{i})\mid i=1,2,\dots,N_{D}\}$, where each $M_{i}$ represents a music signal and $T^{o}_{i}$ is the corresponding original textual description. To optimize model learning from datasets with diverse audio quality and minimize the impact of low-quality audio, we initially assign p-MOS scores to each music track using a model fine-tuned with wav2vec 2.0 (Baevski et al. 2020) on a dataset of vinyl recordings for audio quality assessment, and achieve the corresponding p-MOS set $S=\{s_{1},s_{2},\ldots,s_{n}\}$. These scores facilitate dual-perspective quality control for enhanced granularity and precision.
First, We analyze this p-MOS set $S$ to identify a negative skew normal distribution with mean $\mu$ and variance $\sigma^{2}$. We define text prefixes based on $s$ as follows: prepend “low quality” if $s<\mu-2\sigma$, “medium quality” if $\mu-\sigma\leq m\leq\mu+\sigma$, and “high quality” if $s>\mu+2\sigma$. This information is prepended before processing through the text encoder with cross-attention, enabling the initial separation of quality-related information.
To achieve a more precise awareness and control of waveform quality, we synergize the role of text control with quality embedding. We observed that the distribution of p-MOS in the dataset is approximately normal, which can be shown in Figure 1, allowing us to use the Empirical Rule to segment the data accordingly. Specifically, we define the quantization function $Q:[0,5]\to\{1,2,3,4,5\}$ to map the p-MOS scores to discrete levels based on the distance from the mean $\mu$ in terms of standard deviation $\sigma$:
@@
Quality-aware Masked Diffusion Transformer
In a general patchify phrase with patch size $p_{f}\times p_{l}$ and overlap size $o_{f}\times o_{l}$, patchified token sequence $X=\{x_{1},x_{2},\dots,x_{P}\}\subset\mathbb{R}^{p_{f}\times p_{l}}$ are obtained through spliting the music latent space $\mathcal{M}_{spec}\in\mathbb{R}^{F\times L}$, as described in Section  5. The total number of patches $P$ is given by:
@@
A 2D-Rope position embedding (Su et al. 2024) is added to each patch for better modeling of relative position relationship while a binary mask $\boldsymbol{m}\in\{0,1\}^{P}$ is applied during the training stage, with a variable mask ratio $\gamma$. This results in a subset of $\lfloor\gamma P\rfloor$ patches being masked that $\sum_{i=1}^{P_{N}}m_{i}=\lfloor\gamma P\rfloor$, leaving $P-\lfloor\gamma P\rfloor$ patches unmasked. The subset of masked tokens is invisible in the encoder stage and replaced with trainable mask tokens in the decoder stage following the same strategy utilized in AudioMAE (Huang et al. 2022) and MDT (Gao et al. 2023).
The transformer we use consists of $N$ encoder blocks, $M$ decoder blocks, and an intermediate layer to replace the masked part with trainable parameters. We treat the embedding of the quantized p-MOS score as a prefix token, concatenated with each stage’s music tokens. Let ${X}^{k}=[{x}^{k}_{1},{x}^{k}_{2},\ldots,{x}^{k}_{P}]\in\mathbb{R}^{P\times d}$ represent the output of $k$-th encoder or decoder block, where the initial input of the encoder $X^{0}=z_{t}=\alpha_{t}{z}_{0}+\sqrt{1-\alpha_{t}}\epsilon$, and the final decoder block estimate $X^{N+M}=z_{0}=[{x}_{1},{x}_{2},\ldots,{x}_{P}]$. For $k<N$, indicating the encoder blocks, the sequence transformation focuses only on unmasked tokens:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{(z_{0},q_{vq},y),\epsilon}\left[\left\|\epsilon-D_{\theta}\left(\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,t,q_{vq},{y}\right)\right\|^{2}\right]$ (6)
@@
$D_{\theta}^{\text{High}}(z_{t},t,q_{vq}^{\text{high}},y)$ $=D_{\theta}(z_{t},t,q_{vq}^{\text{high}},y)+$ (7)
$w\left(D_{\theta}(z_{t},t,q_{vq}^{\text{high}},y)-D_{\theta}(z_{t},t,q_{vq}^{\text{low}},\emptyset)\right)$
@@
Music Caption Refinement
@@
Initially, pretrained music caption model (Doh et al. 2023) is employed to re-annotate each music signal $M_{i}$ to $T^{g}_{i}$, shown as $\mathcal{D}_{g}=\{(M_{i},T^{g}_{i})\mid T^{g}_{i}=\mathcal{F}_{\text{cap}}(M_{i}),i=1,2,\dots,N\}$. CLAP text-audio similarity is applied to filter $\mathcal{D}_{g}$ with a threshold of $\rho_{1}$, resulting in
@@
In this context, we meticulously filter out generated captions that do not correspond with their respective audio files. This misalignment may be attributed to inaccuracies within the captioner’s insufficient training. For the filtered data pairs, we opt to retain the use of the original captions.
To ensure that valuable information from the original captions is not overlooked when using only the generated captions, we adapt a fusing stage to combine the original caption and generated pseudo prompt. Firstly, we need to filter out original captions that is useless or inaccurate, formulated as:
@@
The issue can stem from the original data being improperly labeled with terms such as ’speech, car’ from datasets like AudioSet (Gemmeke et al. 2017) and also may be because of desperately missing of the original labels.
@@
$\mathcal{D}_{\text{merge}}=\big{\{}(M_{i},T_{\text{fusion}})\mid\mathcal{S}(T^{o},T^{g})<\rho_{3},$ (10)
$\hskip 25.00003pt(M_{i},T^{o})\in\mathcal{D}^{\text{filter}}_{o},(M_{i},T^{g})\in\mathcal{D}^{\text{filter}}_{g}\big{\}}.$